# IMPACT: Behavioral Intention-aware Multimodal Trajectory Prediction with Adaptive Context Trimming

### v3

@@ 1 Introduction
Figure 1: (a) illustrates the traditional predictor’s context input, while (b) is our integrated approach jointly predicting behavioral intentions, trajectories, and vectorized occupancy. In our approach, the decoder stage is fed only with influential agents and relevant map elements.
With mainstream approaches including [wang2024futurenetlofjointtrajectoryprediction, zhou2024modeseq, shi2024mtr++, Zhou_2023_qc] prioritising generating precise multimodal trajectory predictions, most recent research focused on achieving higher average precisions (mAP) and minimizing deviations (Brier-minFDE). However, humans primarily determine the future behavior of interacting agents not by focusing on specific states alone, but rather by inferring the underlying intention that drives their actions.
@@
Currently, most trajectory prediction models use attention mechanism for querying information from agents and map features. However, indiscriminately attending to all agents and map elements introduces unnecessary complexity and potential causal confusion, as many elements have no direct relevance to the target agent, ultimately hindering correct decoding of the target agent’s future trajectories. To address this, some recent approaches introduce an additional local context-aware refinement module after the decoder stage [zhou2024smartrefinescenarioadaptiverefinementframework, choi2023rpred], but this further complicates the prediction pipeline. Others adopt rule-based heuristic pruning of input lanes [lan2023SEPT] or dynamically collect map elements based on the previous layer prediction results [shi2023MTR, shi2024mtr++, lin2024eda, sun2024rmpyolorobustmotionpredictor]. However, these methods may struggle in complex scenarios or suffer from error propagation. BETOP [liu2024reasoningmultiagentbehavioraltopology] is the first paper explicitly modelling agent interactions through braid theory, but its topological selection algorithm demonstrates critical accuracy deficiencies (see appendix). Thus, an efficient approach to selectively focus on truly critical features (both agents and maps) remains an open challenge. Furthermore, the lack of ground-truth intention labels in current mainstream datasets (e.g., Waymo [waymo_motion_prediction], Argoverse [argoverse, argoverse2]) is limiting the development of accurate behavioral intention-aware models.
@@
To holistically address the aforementioned challenges, we propose our approach, IMPACT. Specifically, we introduce an agent-wise multimodal behavioral intention module before the trajectory decoder, while sharing the same scenario information extraction module (i.e., the same encoder) used by the trajectory predictor. This design reduces computational overhead and mitigates information loss. The agent-wise behavioral intention module outputs a one-hot intention vector for each vehicle, where each element in the vector represents the probability of a specific intention relative to the ego vehicle (e.g., overtaking, yielding). Based on these probability vectors, we apply a utility function to assign an “interaction score” to each vehicle, selecting only those with high interaction probabilities as inputs to the decoder. In parallel, we propose a vectorized occupancy prediction module. This module predicts the probability that each vectorized map element will be occupied by the ego vehicle in the future. Similar to the agent selection process, we only feed the map polylines with high occupancy probabilities into the decoder. By integrating both modules in this manner, our method effectively reduces redundant interactions while preserving crucial information for accurate and robust trajectory prediction. We also propose an automatic labeling algorithm to generate high-quality ground truth labels for behavioral intentions on the Waymo and Argoverse 1&2 dataset.
@@
We evaluated IMPACT on both the Waymo Motion Prediction Dataset (for marginal prediction) and the Waymo Interactive Prediction Dataset (for interaction prediction). Experimental results demonstrate that our method achieves the best performance among all LiDAR-free methods on the Waymo Motion Prediction Dataset, surpassed only by MTRV3 [shi2024mtrv3], which leverages additional LiDAR data. Moreover, IMPACT ranks first overall on the Waymo Interactive Prediction Dataset, surpassing the previous state-of-the-art ensemble model [liu2024reasoningmultiagentbehavioraltopology] using only a single model. To further demonstrate the practical applicability of our approach, we successfully trained IMPACT on a proprietary dataset and deployed it effectively on a real vehicle.
@@ 2.1 Motion Prediction
Early approaches [chai2019multipathmultipleprobabilisticanchor, gilles2021homeheatmapoutputfuture, phanminh2020covernetmultimodalbehaviorprediction] relied on Bird’s-Eye-View (BEV) rasterized representations processed through CNNs, but struggled to preserve geometric fidelity and semantic relationships. This limitation spurred vectorized paradigms like VectorNet [gao2020vectornetencodinghdmaps], which encoded agents and map elements as polylines. Subsequent graph-based architectures [deo2021multimodaltrajectorypredictionconditioned, JiaHDGT] further formalized relational dependencies through graph neural networks. The field then diverged into anchor-oriented strategies [zhou2024smartrefinescenarioadaptiverefinementframework, MultiPath++, gu2021densetnt, zhao2020tnt] leveraging HD map anchors for physically constrained prediction, followed by agent-centric frameworks [Hivt, shi2023MTR]. But it requires re-normalizing and re-encoding during online inference. Query-centric approaches[zhou2023qcnext, Zhou_2023_qc, zhang2024simple, shi2024mtr++, futurelof] addressed this by encoding each vehicle as a query, leveraging attention mechanisms to enhance prediction performance and improve multi-agent coherence through unified spatio-temporal encoding. Recent research in motion forecasting has witnessed significant advancements across multiple frontiers, driving improvements in generalization, consistency, and predictive accuracy. Self-supervised approaches [bhattacharyya2022ssllane, forecastmae, lan2023SEPT] enhance model generalization by leveraging reconstruction pretext tasks, while temporal consistency methods [zhang2024demodecouplingmotionforecasting, tang2024hpnet, realmotion] reinforce scene coherence through continuous-time modeling. At the same time, architectures like RMP-YOLO [sun2024rmpyolorobustmotionpredictor, POP_RAL] tackle the challenge of partial observations by imputing missing trajectory segments. Furthermore, ModeSeq [zhou2024modeseq] introduces a paradigm shift by framing motion prediction as a GPT-style next-token prediction task.
However, current methods tend to emphasize the precise generation of multimodal trajectory predictions at the expense of behavioral intention prediction[fang2023behavioralintentionpredictiondriving] (BIP). This narrow focus often overlooks the latent decision-making processes that drive observable maneuvers, resulting in models that struggle to interpret complex social interactions or anticipate nuanced changes in driving behavior. In real-world scenarios, the inability to infer intentions such as overtake or yield decisions can lead to suboptimal planning and reduced overall robustness. The most recent work, BeTOP [liu2024reasoningmultiagentbehavioraltopology], attempts to explicitly model behavioral interactions via braid theory by splitting agents into “interactive” and “non-interactive” based on whether their future trajectories form an intertwined braid. However, in most scenarios (see Appendix), the generated braids appear unrealistic because the approach only considers lateral and temporal dimensions. Moreover, relying on a binary behavioral label oversimplifies complex traffic scenarios, where interactions can be far more nuanced.
@@ 2.2 Context-aware Pruning
During the decoder stage, current works often relies on attention mechanisms [vaswani2023attentionneed] to query agent-map information. However, using global attention over all entities incurs $O(n^{2})$ complexity and can introduce causal confusion, in contrast to human drivers, who primarily focus on goal-critical paths and potentially interactive vehicles [fang2023behavioralintentionpredictiondriving]. Some approaches address this via prior rule-based pruning (e.g., SEPT [lan2023SEPT]), which may fail for corner case. Meanwhile, MTR-series[shi2023MTR, shi2024mtr++, sun2024controlmtr, lin2023eda, mu2024mostmultimodalityscenetokenization, gan2024mgtrmultigranulartransformermotion, sun2024rmpyolorobustmotionpredictor, llm_augmented_mtr] methods and R-pred[choi2023rpred] use last layer predictions as priors to dynamically collect nearby polylines, yet suffers from error propogation. BeTOP[liu2024reasoningmultiagentbehavioraltopology], which relies on braid theory to select interactive vehicles, can yield an unreasonably chosen set of vehicles in practice. To tackle these issues, our IMPACT framework introduces a symmetric dual context-filtering approach that leverage predicted behavioral interaction probabilities and vectorized occupancy to retain only influential vehicles and relevant map elements.
@@ 3.3.1 Spatial Temporal Encoding
We utilize RMP-YOLO’s encoder[sun2024rmpyolorobustmotionpredictor] as a backbone network (see Figure 2) to do feature fusion. To integrate heterogeneous input modalities, we employ a cascaded Multi-Context Gating (MCG) mechanism inspired by [MultiPath++]. The MCG modules sequentially fuse pairs of modalities from a candidate set of three. The output of each MCG stage serves as input to the subsequent stage, enabling hierarchical feature interaction:
@@ 3.3.3 Multimodal Behavioral Intention Prediction
Each vector element $h$ represents a probability distribution over four intention categories: yielding, overtaking, ignored, and nearby. To make the decoding process more focused, we first compute an overall interaction score from the predicted distribution $\hat{\mathrm{H}}$ using a utility function $\psi$, producing $\psi(\hat{\mathrm{H}^{i}})\in\mathbb{R}^{K\times N_{a}}$. We then select the top-$m$ highest-scoring agents for downstream trajectory decoding. Therefore, for each modality, we choose $m$ most revelant agents $A_{5}\in\mathbb{R}^{m\times D}$, where $m\ll N_{a}$. This filtering step refines the decoder’s input, concentrating on the most influential interactions while improving prediction accuracy. The ground-truth label of behavioral intention $\mathrm{H}^{*}$ is derived from an auto-labeled data preprocessing process (see Appendix).
@@ 3.3.4 Multimodal Vectorized Occupancy Prediction
This vectorized approach ensures both efficiency and scalability while maintaining alignment with the overall vectorized representation of the scene. Among the $N_{l}$ polylines’s multimodal occupancy probabilities $\varphi(\hat{\mathrm{O}^{i}})\in\mathbb{R}^{K\times N_{l}}$, we select the top-$n$ with the highest predicted probabilities in each modality to form $L_{5}\in\mathbb{R}^{n\times D}$, where $n\ll N_{l}$. These top-ranked polylines serve as focused inputs for the subsequent trajectory decoder. The ground-truth occupancy label $\mathrm{O}^{*}$ is also derived from an auto-labeled data preprocessing process (see Appendix).
@@ 3.3.5 Trajectory Decoder
We adopt a multi-layer MTR-style[shi2023MTR] trajectory decoder. At each layer $i$, self-attention is applied to the query content $Q^{i}\in\mathbb{R}^{K\times D}$ across the $K$ motion modes, enabling information exchange among different future modalities. Subsequently, for each modality, two cross-attention modules integrate features from the filtered agent tokens $A_{5}$ and map tokens $L_{5}$. Finally, the target agent feature (replicated $K$ times) is concatenated with the cross-attended query features, and passed through a regression head to generate a set of Gaussian Mixture Model (GMM) parameters at each future timestep: $\Bigl\{\bigl(\mu_{x}^{k},\mu_{y}^{k},\sigma_{x}^{k},\sigma_{y}^{k},\rho^{k}\bigr)\Bigr\}_{k=1}^{K},$ where $\bigl(\mu_{x}^{k},\mu_{y}^{k},\sigma_{x}^{k},\sigma_{y}^{k},\rho^{k}\bigr)$ parameterizes the $k$-th Gaussian component. In addition, a classification head outputs the confidence scores $S\in\mathbb{R}^{K}$ corresponding to each motion mode. This multimodal representation captures the inherent uncertainties of agent trajectories.
@@ 3.4 Training Loss
where $\lambda_{\mathrm{1}},\lambda_{\mathrm{2}},\lambda_{\mathrm{3}}$, and $\lambda_{\mathrm{4}}$ are weighting factors balancing the contributions of behavioral intention prediction, vectorized occupancy prediction, trajectory prediction, and mode classification, respectively. Specifically, $\mathcal{L}_{\mathrm{Int}}$ is calculated using the multi-class Focal Loss, $\mathcal{L}_{\mathrm{Occ}}$ is based on the binary Focal Loss, $\mathcal{L}_{\mathrm{Traj}}$ is derived from the GMM loss, and $\mathcal{L}_{\mathrm{Score}}$ is computed with Binary Cross-Entropy. During training, the winner-take-all strategy is applied for $\mathcal{L}_{\mathrm{Int}}$, $\mathcal{L}_{\mathrm{Occ}}$, and $\mathcal{L}_{\mathrm{Traj}}$, ensuring that only the modality closest to the ground-truth trajectory is used to compute these losses. Please check appendix for more details.
@@ 4.1 Experimental Setup
Our experiments are conducted on one of the most challenging prediction datasets, Waymo Open Motion Dataset (WOMD)[waymo_motion_prediction]. This large-scale dataset comprises 486,995 training clips, 44,097 validation clips, and 44920 testing clips. Each clip contains 10 timesteps of historical agent states, 1 current timestep, and 80 future timesteps at a sampling frequency of 10 Hz, along with HD map information. We evaluate our method on both core WOMD tasks: the marginal motion and interactive motion prediction tasks. Following the standard evaluation protocol, we adopt metrics including Soft mAP, mAP, minADE, minFDE, Miss Rate, and Overlap Rate, and Soft mAP is the main metric.
@@
We employ AdamW optimizer [loshchilov2019decoupledweightdecayregularization] for training, conducting experiments on a cluster of 8 NVIDIA A800 GPUs with a total batch size of 80. The learning rate is initialized as $1\times 10^{-4}$ and begins step decay starting at epoch 22, halving every two epochs. The model undergoes 30 epochs. More details can be referred to Appendix.
@@ 4.2 Leaderboard Performance
Figure 4 illustrates the detailed qualitative results for both marginal and joint predictions using our proposed method. LABEL:fig:figure-sampling-space shows our qualitative marginal and joint prediction results. For additional qualitative results, please refer to Appendix.

### v4

@@ 1 Introduction
Figure 1: (a) illustrates the traditional context input, while (b) is our integrated approach jointly predicting behavioral intentions, trajectories, and vectorized occupancy. In our approach, the decoder stage is fed only with influential agents and relevant map elements.
With mainstream approaches including [37, 47, 28, 46] prioritising generating precise multimodal trajectory predictions, most recent research focused on achieving higher average precisions (mAP) and minimizing deviations (Brier-minFDE). However, humans primarily determine the future behavior of interacting agents not by focusing on specific states alone, but rather by inferring the underlying intention that drives their actions.
@@
Currently, most trajectory prediction models use attention mechanism for querying information from agents and map features. However, indiscriminately attending to all agents and map elements introduces unnecessary complexity and potential causal confusion, as many elements have no direct relevance to the target agent, ultimately hindering correct decoding of the target agent’s future trajectories. To address this, some recent approaches introduce an additional local context-aware refinement module after the decoder stage [45, 6], but this further complicates the prediction pipeline. Others adopt rule-based heuristic pruning of input lanes [17] or dynamically collect map elements based on the previous layer prediction results [27, 28, 19, 30]. However, these methods may struggle in complex scenarios or suffer from error propagation. BETOP [21] is the first paper explicitly modelling agent interactions through braid theory, but its topological selection algorithm demonstrates critical accuracy deficiencies (see appendix Figure 9). Thus, an efficient approach to selectively focus on truly critical features (both agents and maps) remains an open challenge. Furthermore, the lack of ground-truth intention labels in current mainstream datasets (e.g., Waymo [39], Argoverse [1, 2]) is limiting the development of accurate behavioral intention-aware models.
@@
To holistically address the aforementioned challenges, we propose our approach, IMPACT. Specifically, we introduce an agent-wise multimodal behavioral intention module before the trajectory decoder, while sharing the same scenario information extraction module (i.e., the same encoder) used by the trajectory predictor. This design reduces computational overhead and mitigates information loss. The agent-wise behavioral intention module outputs a one-hot intention vector for each vehicle, where each element in the vector represents the probability of a specific intention relative to the target vehicle (e.g., overtaking, yielding). Based on these probability vectors, we apply a utility function to assign an “interaction score” to each vehicle, selecting only those with high interaction probabilities as inputs to the decoder. In parallel, we propose a vectorized occupancy prediction module. This module predicts the probability that each vectorized map element will be occupied by the target vehicle in the future. Similar to the agent selection process, we only feed the map polylines with high occupancy probabilities into the decoder. By integrating both modules in this manner, our method effectively reduces redundant interactions while preserving crucial information for accurate and robust trajectory prediction. We also propose an automatic labeling algorithm to generate high-quality ground truth labels for behavioral intentions on the Waymo and Argoverse 1&2 dataset.
@@
We evaluated IMPACT on both the Waymo Motion Prediction Dataset (for marginal prediction) and the Waymo Interactive Prediction Dataset (for interaction prediction). Experimental results demonstrate that our method achieves the best performance among all LiDAR-free methods on the Waymo Motion Prediction Dataset, surpassed only by MTRV3 [26], which leverages additional LiDAR data. Moreover, IMPACT ranks first overall on the Waymo Interactive Prediction Dataset, surpassing the previous state-of-the-art ensemble model [21] using only a single model. To further demonstrate the practical applicability of our approach, we successfully trained IMPACT on a proprietary dataset and deployed it effectively on a real vehicle.
@@ 2.1 Motion Prediction
Early approaches [4, 12, 25] relied on Bird’s-Eye-View (BEV) rasterized representations processed through CNNs, but struggled to preserve geometric fidelity and semantic relationships. This limitation spurred vectorized paradigms like VectorNet [11], which encoded agents and map elements as polylines. Subsequent graph-based architectures [7, 15] further formalized relational dependencies through graph neural networks. The field then diverged into anchor-oriented strategies [45, 34, 13, 43] leveraging HD map anchors for physically constrained prediction, followed by agent-centric frameworks [49, 27]. But it requires re-normalizing and re-encoding during online inference. Query-centric approaches[48, 46, 41, 28, 36] addressed this by encoding each vehicle as a query, leveraging attention mechanisms to enhance prediction performance and improve multi-agent coherence through unified spatio-temporal encoding. Recent research in motion forecasting has witnessed significant advancements across multiple frontiers, driving improvements in generalization, consistency, and predictive accuracy. Self-supervised approaches [3, 5, 17] enhance model generalization by leveraging reconstruction pretext tasks, while temporal consistency methods [40, 33, 29] reinforce scene coherence through continuous-time modeling. At the same time, architectures like RMP-YOLO [30, 38] tackle the challenge of partial observations by imputing missing trajectory segments. Furthermore, ModeSeq [47] introduces a paradigm shift by framing motion prediction as a GPT-style next-token prediction task.
However, current methods tend to emphasize the precise generation of multimodal trajectory predictions at the expense of behavioral intention prediction[8] (BIP). This narrow focus often overlooks the latent decision-making processes that drive observable maneuvers, resulting in models that struggle to interpret complex social interactions or anticipate nuanced changes in driving behavior. In real-world scenarios, the inability to infer intentions such as overtake or yield decisions can lead to suboptimal planning and reduced overall robustness. The most recent work, BeTOP [21], attempts to explicitly model behavioral interactions via braid theory by splitting agents into “interactive” and “non-interactive” based on whether their future trajectories form an intertwined braid. However, in most scenarios (see Appendix Figure 9), the generated braids appear unrealistic because the approach only considers lateral and temporal dimensions. Moreover, relying on a binary behavioral label oversimplifies complex traffic scenarios, where interactions can be far more nuanced.
@@ 2.2 Context-aware Pruning
During the decoder stage, current works often relies on attention mechanisms [35] to query agent-map information. However, using global attention over all entities incurs $O(n^{2})$ complexity and can introduce causal confusion, in contrast to human drivers, who primarily focus on goal-critical paths and potentially interactive vehicles [8]. Some approaches address this via prior rule-based pruning (e.g., SEPT [17]), which may fail for corner case. Meanwhile, MTR-series[27, 28, 31, 18, 23, 10, 30, 44] methods and R-pred[6] use last layer predictions as priors to dynamically collect nearby polylines, yet suffers from error propogation. BeTOP[21], which relies on braid theory to select interactive vehicles, can yield an unreasonably chosen set of vehicles in practice. To tackle these issues, our IMPACT framework introduces a symmetric dual context-filtering approach that leverage predicted behavioral interaction probabilities and vectorized occupancy to retain only influential vehicles and relevant map elements.
@@ 3.3.1 Spatial Temporal Encoding
We utilize RMP-YOLO’s encoder[30] as a backbone network (see Figure 2) to do feature fusion. To integrate heterogeneous input modalities, we employ a cascaded Multi-Context Gating (MCG) mechanism inspired by [34]. The MCG modules sequentially fuse pairs of modalities from a candidate set of three. The output of each MCG stage serves as input to the subsequent stage, enabling hierarchical feature interaction:
@@ 3.3.3 Multimodal Behavioral Intention Prediction
Each vector element $h$ represents a probability distribution over four intention categories: yielding, overtaking, ignored, and nearby. To make the decoding process more focused, we first compute an overall interaction score from the predicted distribution $\hat{\mathrm{H}}$ using a utility function $\psi$, producing $\psi(\hat{\mathrm{H}^{i}})\in\mathbb{R}^{K\times N_{a}}$. We then select the top-$m$ highest-scoring agents for downstream trajectory decoding. Therefore, for each modality, we choose $m$ most revelant agents $A_{5}\in\mathbb{R}^{m\times D}$, where $m\ll N_{a}$. This filtering step refines the decoder’s input, concentrating on the most influential interactions while improving prediction accuracy. The ground-truth label of behavioral intention $\mathrm{H}^{*}$ is derived from an auto-labeled data preprocessing process (see Appendix Algorithm 2).
@@ 3.3.4 Multimodal Vectorized Occupancy Prediction
This vectorized approach ensures both efficiency and scalability while maintaining alignment with the overall vectorized representation of the scene. Among the $N_{l}$ polylines’s multimodal occupancy probabilities $\varphi(\hat{\mathrm{O}^{i}})\in\mathbb{R}^{K\times N_{l}}$, we select the top-$n$ with the highest predicted probabilities in each modality to form $L_{5}\in\mathbb{R}^{n\times D}$, where $n\ll N_{l}$. These top-ranked polylines serve as focused inputs for the subsequent trajectory decoder. The ground-truth occupancy label $\mathrm{O}^{*}$ is also derived from an auto-labeled data preprocessing process (see Appendix Algorithm 1).
@@ 3.3.5 Trajectory Decoder
We adopt a multi-layer MTR-style[27] trajectory decoder. At each layer $i$, self-attention is applied to the query content $Q^{i}\in\mathbb{R}^{K\times D}$ across the $K$ motion modes, enabling information exchange among different future modalities. Subsequently, for each modality, two cross-attention modules integrate features from the filtered agent tokens $A_{5}$ and map tokens $L_{5}$. Finally, the target agent feature (replicated $K$ times) is concatenated with the cross-attended query features, and passed through a regression head to generate a set of Gaussian Mixture Model (GMM) parameters at each future timestep: $\Bigl\{\bigl(\mu_{x}^{k},\mu_{y}^{k},\sigma_{x}^{k},\sigma_{y}^{k},\rho^{k}\bigr)\Bigr\}_{k=1}^{K},$ where $\bigl(\mu_{x}^{k},\mu_{y}^{k},\sigma_{x}^{k},\sigma_{y}^{k},\rho^{k}\bigr)$ parameterizes the $k$-th Gaussian component. In addition, a classification head outputs the confidence scores $S\in\mathbb{R}^{K}$ corresponding to each motion mode. This multimodal representation captures the inherent uncertainties of agent trajectories.
@@ 3.4 Training Loss
where $\lambda_{\mathrm{1}},\lambda_{\mathrm{2}},\lambda_{\mathrm{3}}$, and $\lambda_{\mathrm{4}}$ are weighting factors balancing the contributions of behavioral intention prediction, vectorized occupancy prediction, trajectory prediction, and mode classification, respectively. Specifically, $\mathcal{L}_{\mathrm{Int}}$ is calculated using the multi-class Focal Loss, $\mathcal{L}_{\mathrm{Occ}}$ is based on the binary Focal Loss, $\mathcal{L}_{\mathrm{Traj}}$ is derived from the GMM loss, and $\mathcal{L}_{\mathrm{Score}}$ is computed with Binary Cross-Entropy. During training, the winner-take-all strategy is applied for $\mathcal{L}_{\mathrm{Int}}$, $\mathcal{L}_{\mathrm{Occ}}$, and $\mathcal{L}_{\mathrm{Traj}}$, ensuring that only the modality closest to the ground-truth trajectory is used to compute these losses. Please check appendix (section 2) for more details.
@@ 4.1 Experimental Setup
Our experiments are conducted on one of the most challenging prediction datasets, Waymo Open Motion Dataset (WOMD)[39]. This large-scale dataset comprises 486,995 training clips, 44,097 validation clips, and 44920 testing clips. Each clip contains 10 timesteps of historical agent states, 1 current timestep, and 80 future timesteps at a sampling frequency of 10 Hz, along with HD map information. We evaluate our method on both core WOMD tasks: the marginal motion and interactive motion prediction tasks. Following the standard evaluation protocol, we adopt metrics including Soft mAP, mAP, minADE, minFDE, Miss Rate, and Overlap Rate, and Soft mAP is the main metric.
@@
We employ AdamW optimizer [22] for training, conducting experiments on a cluster of 8 NVIDIA A800 GPUs with a total batch size of 80. The learning rate is initialized as $1\times 10^{-4}$ and begins step decay starting at epoch 22, halving every two epochs. The model undergoes 30 epochs. More details can be referred to Appendix section 4.
@@ 4.2 Leaderboard Performance
Figure 5 illustrates the detailed qualitative results for both marginal and joint predictions using our proposed method. For additional qualitative results, please refer to section 7 in Appendix.
RMP-YOLO(Ensemble)[30] 0.4737 0.4531 0.5564 1.1188 0.1084 ${0.1259}^{*}$
ModeSeq(Ensemble)[47] 0.4737 0.4665 0.5680 1.1766 0.1204 0.1275
BeTop[21] 0.4698 0.4587 0.5716 1.1668 0.1183 0.1272
MGTR[10] 0.4599 0.4505 0.5918 1.2135 0.1298 0.1275
EDA[19] 0.4596 0.4487 0.5718 1.1702 0.1169 0.1266
MTR++[28] 0.4410 0.4329 0.5906 1.1939 0.1298 0.1281
MTR[27] 0.4403 0.4249 0.5964 1.2039 0.1312 0.1274
HPTR[42] 0.3968 0.3904 0.5565 1.1393 0.1434 0.1366
HDGT[15] 0.3709 0.3577 0.5933 1.2055 0.1511 0.1557
DenseTNT[13] - 0.3281 1.0387 1.5514 0.1573 0.1779
SceneTransformer[24] - 0.2788 0.6117 1.2116 0.1564 0.1473
Ours (Ensemble) 0.4801 ${0.4598}^{*}$ 0.5563 1.1295 0.1087 0.1258
Ours (Single) 0.4721 0.4609 ${0.5641}^{*}$ ${1.1540}^{*}$ ${0.1143}^{*}$ 0.1255
BeTop-ens [21] 0.2573 0.2511 0.9779 2.2805 0.4376 $0.1688^{*}$
BeTop [21] $0.2466^{*}$ $0.2412^{*}$ 0.9744 2.2744 0.4355 0.1696
MTR++ [28] 0.2368 0.2326 $\underline{0.8975}$ $\underline{1.9509}$ 0.4143 0.1665
MTR [27] 0.2078 0.2037 $0.9181^{*}$ 2.0633 0.4411 0.1717
MotionDiffuser [16] 0.2047 0.1952 0.8642 1.9482 $\underline{0.4300}$ 0.2004
GameFormer [14] 0.1982 0.1923 0.9721 2.2146 0.4933 0.2022
BIFF [50] 0.1383 0.1229 0.9388 $2.0443^{*}$ 0.4620 0.2172
DenseTNT [13] - 0.1647 1.1417 2.4904 0.5350 0.2309
M2I [32] - 0.1239 1.3506 2.8325 0.5538 0.2757
SceneTransformer [24] - 0.1192 0.9774 2.1892 0.4942 0.2067
Ours (Single) 0.2718 0.2659 0.9738 2.2734 $0.4316^{*}$ 0.1684
4.3 Ablation Study
Different behavior modeling and map interaction mechanisms.
Table 3 presents an ablation study evaluating various agent-pruning and map-pruning strategies on the validation set. The first row (All + Dynamic) queries all agents for the decoder while employing dynamic map selection as in MTR [27], serving as a baseline. Introducing agent selection via Braid Theory [21] (Row 2) enhances Soft mAP to 0.4602. Replacing Braid Theory with our two-class filtering—where Yielding, Overtaking, and Nearby behaviors are treated as interactive while Ignore is considered non-interactive (Row 3)—further refines minADE to 0.5693 while maintaining a comparable Miss Rate. Notably, adopting our four-class YOIN approach (Row 4) elevates Soft mAP to 0.4728, demonstrating its ability to better capture nuanced interactions. Lastly, incorporating vectorized occupancy into YOIN (Row 5) further improves performance (Soft mAP = 0.4758), highlighting the benefit of explicit map filtering in isolating relevant polylines and refining trajectory predictions. These results validate the effectiveness of our integrated behavior modeling and map selection framework.
Agent Map Metrics
All w/Braid Theory[21] w/Inter. w/YOIN w/Dynamic[27] w/Vect. Occ. Soft mAP mAP minADE Miss Rate Inference Time (ms/scenario) Params (M)
0.4582 0.4414 0.5718 0.1203 14.67 69.469
0.4602 0.4507 0.5745 0.1187 23.92 48.938
0.4663 0.4549 0.5693 0.1187 23.92 48.938
0.4728 0.4613 0.5680 0.1163 23.92 48.941
0.4758 0.4646 0.5652 0.1173 18.27 48.944
Marginal Prediction Performance.
Table 1 shows that our method achieves state-of-the-art (SOTA) performance on the Waymo 2024 motion prediction benchmark, outperforming all existing LiDAR-free approaches in the primary metric, Soft mAP, as well as in minADE and Overlap Rate. In the remaining metrics (mAP, minFDE, and Miss Rate), our method ranks second. Here, ’Ensemble’ denotes the model-ensemble technique described in [27] for performance boosting, while ’Single’ indicates evaluation with a single model.